ax/2b-c/3d+x=1 solve for x
Answer by stanbon(75887)

ax/2b-c/3d+x=1 solve for x
-------
[ax/2b] - [c/3d] + x - 1
[a/2b + 1]x = 1 + [c/3d]
[(a+2b)/2b]x = [(3d+c)/3d

x = [(3d+c)/3d] * {2b/(a+2b)]
========================
Cheers,
Stan H.